### v1

# ALCM: Autonomous LLM-Augmented Causal Discovery Framework

Abstract
To perform effective causal inference in high-dimensional datasets, initiating the process with causal discovery is imperative, wherein a causal graph is generated based on observational data. However, obtaining a complete and accurate causal graph poses a formidable challenge, recognized as an NP-hard problem. Recently, the advent of Large Language Models (LLMs) has ushered in a new era, indicating their emergent capabilities and widespread applicability in facilitating causal reasoning across diverse domains, such as medicine, finance, and science. The expansive knowledge base of LLMs holds the potential to elevate the field of causal reasoning by offering interpretability, making inferences, generalizability, and uncovering novel causal structures. In this paper, we introduce a new framework, named Autonomous LLM-Augmented Causal Discovery Framework (ALCM), to synergize data-driven causal discovery algorithms and LLMs, automating the generation of a more resilient, accurate, and explicable causal graph. The ALCM consists of three integral components: causal structure learning, causal wrapper, and LLM-driven causal refiner. These components autonomously collaborate within a dynamic environment to address causal discovery questions and deliver plausible causal graphs. We evaluate the ALCM framework by implementing two demonstrations on seven well-known datasets. Experimental results demonstrate that ALCM outperforms existing LLM methods and conventional data-driven causal reasoning mechanisms. This study not only shows the effectiveness of the ALCM but also underscores new research directions in leveraging the causal reasoning capabilities of LLMs.
Keywords: Large Language Models; Causal Reasoning; Causal Graph; Causal Discovery;
1 Introduction
The process of causal discovery, essential in various domains and scientific discoveries, seeks to reveal complex causal relationships in observational data [31, 32, 13]. For instance, in healthcare, this process is crucial and instrumental for pinpointing disease etiologies, devising effective interventions, and prevention strategies [48]. Subsequently, causal inference allows for the quantification of the influence exerted by different variables on one another, once a causal structure is identified. This phase, often referred to as causal estimation, relies on the construction of a preliminary causal graph, which, despite its theoretical significance, poses considerable practical challenges, demanding substantial domain-specific expertise. In fact, studies using real-world datasets demonstrate that inferring causal graphs–which is the focus of this paper–from data is still a complex challenge in practical applications [34, 42, 6]. Causal discovery and causal inference, as highlighted in seminal works by Pearl and others [31, 32, 23, 13], are two key components of causal reasoning to address causal questions in diverse fields.
Within the literature, numerous studies have contributed significantly to the development of a variety of efficient causal discovery algorithms aimed at uncovering the underlying causal structure from observational data. This body of research can be broadly categorized into two main groups: conventional data-driven causal discovery algorithms and those based on LLMs. Conventional causal discovery algorithms focus on learning the causal graph from samples of the joint probability distribution of observational data. They utilize various statistical techniques, including conditional independence tests, machine learning approaches, generative models, deep learning methodologies, and reinforcement learning strategies [35] to understand the joint distribution of observed variables and extract the causal connections among them. Subsequently, these algorithms assess how well the candidate causal graph aligns with the data [48, 14, 13].
Conventional causal discovery algorithms, despite being designed to be powerful and scalable, face several challenges. These include a heavy dependence on domain experts [12], who are often limited and inconsistent, and the issues of data bias, imbalance, and inadequacy which affect the accuracy of capturing true probability distributions [6]. Additionally, the use of static data can compromise model accuracy in dynamic environments, and the task of fully determining edge orientations is hindered by the presence of multiple equivalent Directed Acyclic Graphs (DAGs) [6, 35], which exponentially increase with the number of nodes [50], leading to inaccuracies and unreliability in the estimated causal graphs.
Recent advancements in Large Language Models (LLMs) have significantly impacted artificial intelligence, exhibiting notable reasoning capabilities [20, 44, 9, 22, 4]. These achievements stem from the extensive data used for training LLMs, essential for effective causal reasoning [20, 9]. However, current LLM-based causal reasoning research, mainly focusing on pairwise analysis, faces scalability issues as it struggles with the complexity of full causal graph construction and handling large datasets [43, 20, 6, 7, 30]. These models often fall short in accurately and efficiently inferring comprehensive causal relationships, especially in large-scale settings [7, 6, 23, 18]. Despite some efforts to integrate LLMs with causal discovery processes [43, 7, 39], challenges remain due to inherent limitations and the complexity of causal inference. A synergistic approach combining LLMs with other methods may provide a more nuanced and complete understanding of causal mechanisms and address these challenges effectively.
In this paper, we present an LLM-powered causal discovery framework–ALCM: a multi-component Autonomous LLM-Augmented Causal Discovery Framework. ALCM proposes a synergized reasoning method and entails three components: causal structure learning, causal wrapper, and LLM-driven refiner components to generate more accurate and robust causal graphs. ALCM is engineered to autonomously untangle causal structures by deciphering those causal relations embedded in observational data. ALCM capitalizes on observed data, data-driven causal reasoning algorithms, and the implicit knowledge embedded in LLMs to optimize and streamline the entire causal reasoning process. This approach aims to establish a more robust, applicable, and reliable foundation for causal reasoning and estimation as well. We conduct a comprehensive performance evaluation of ALCM, employing LLMs and assessing their capabilities on widely recognized benchmarks [36, 42]. We compare our framework with conventional causal discovery algorithms and LLMs prompting. Furthermore, we implement an automatic pipeline for making the causal discovery an automatic task.
2 Background and Related Work
In this section, we outline the existing research on causal structure learning within the literature, delineating it into two primary groups: 1) Conventional data-driven causal discovery algorithms; and 2) Using LLMs for causal discovery.
1) Conventional data-driven causal discovery algorithms: conventional data-driven causal discovery algorithms are broadly classified into five categories as follows:
  • Score-Based Algorithms: They operate on scores and engage in a comprehensive exploration of the entire space of potential Directed Acyclic Graphs (DAGs) to identify the most suitable graph for explaining the underlying data. Typically, such score-based approaches consist of two integral components: (i) a systematic search strategy tasked with navigating through the potential search states or the space of candidate graphs, denoted as G’, and (ii) a score function designed to evaluate the viability of these candidate causal graphs. The synergy between the search strategy and the score function is instrumental in optimizing the exploration of all conceivable DAGs. A widely employed score function in the selection of causal models is the Bayesian Information Criterion (BIC) [14]. Some examples of score-based algorithms are Greedy Equivalence Search (GES) [11], Fast Greedy Search (FGS) [33], and A* Search [47].
  • Constraint-Based Algorithms: This category, exemplified by Peter-Clark (PC) [38] algorithm, employs conditional independence (CI) tests to reveal the graph’s skeleton and v-structures, ultimately returning the Directed Acyclic Graph (DAG) of the functional causal model while considering v-structures and doing edge-orientations [14]. Other constraint-bsaed algorithms are like Fast Causal Inference (FCI), Anytime FCI, RFCI, PC-stable, and so forth.
  • Hybrid Algorithms: Hybrid approaches are founded on the integration of various causal discovery methods, combining constraint-based, score-based, Functional Causal Model (FCM)-based, gradient-based, and other techniques. This amalgamation reflects a comprehensive strategy that leverages the strengths of different methodologies to enhance the robustness and effectiveness of causal discovery in complex systems. Max-Min Hill Climbing (MMHC) [40]–belonging to this category–stands out as a hybrid causal discovery technique that seamlessly integrates principles from both score-based and constraint-based algorithms. This hybrid approach combines the advantages of scoring methods and constraint-based strategies, offering a comprehensive and effective framework for uncovering causal relationships in complex systems.
  • Function-Based Algorithms: Approaches grounded in Functional Causal Models (FCM) delineate the causal connections between variables within a defined functional structure. In FCMs, variables are expressed as functions of their direct causes (parents), augmented by an independent noise term denoted as E. The distinguishing feature of FCM-based methodologies lies in their capacity to differentiate between various Directed Acyclic Graphs (DAGs) within the same equivalence class. This discrimination is achieved by introducing supplementary assumptions concerning data distributions and/or function classes. Several notable FCM-based causal discovery methodologies are introduced, including Linear Non-Gaussian Acyclic Model (LiNGAM) [37] and Structural Agnostic Modeling (SAM) [19]. SAM employs an adversarial learning methodology for causal graph identification. Specifically, SAM utilizes Generative Adversarial Neural Networks (GANs) to seek a Functional Causal Model (FCM) while ensuring the detection of sparse causal graphs through the incorporation of appropriate regularization terms. The optimization process involves a learning criterion that integrates distribution estimation, sparsity considerations, and acyclicity constraints. This holistic criterion facilitates end-to-end optimization of both the graph structure and associated parameters, accomplished through stochastic gradient descent.
    The previous three-mentioned categories may be limited to the Markov equivalence class, posing constraints. Function-based algorithms like LiNGAM [44] aim to uniquely identify causal DAGs by exploiting data generative process asymmetries or causal footprints.
  • Optimization-Based Algorithms: Recent investigations in causal discovery have approached the structure learning problem by casting it as a continuous optimization task, employing the least squares objective and an algebraic representation of Directed Acyclic Graphs (DAGs). Notably, this transformation converts the combinatorial nature of the structure learning problem into a continuous framework, and solutions are obtained through the application of gradient-based optimization techniques. These methods exploit the gradients of an objective function concerning the parameterization of a DAG matrix to achieve effective structure learning. NOTEARS [50] is among the causal discovery algorithms that formulate the structure learning problem as a purely continuous constrained optimization task.
2) Using LLM for causal discovery task: Leveraging recent advancements in LLMs and Natural Language Processing (NLP) presents an opportunity to offer enhanced capabilities in capturing causal concepts and relations while handling large-scale datasets more effectively [26, 10, 27]. This proficiency is rooted in the extensive training LLMs undergo on vast, high-quality datasets [18]. LLMs possess the ability to establish a comprehensive knowledge base across diverse domains, facilitating language understanding, ensuring generalizability, automating the causal reasoning pipeline, and enabling plausible reasoning. In this regard, the second group, namely using LLMs for causal discovery, is introduced. This group is classified into three major groups as follows:
  • Fine-tuning: This category mainly focuses on fine-tuning LLMs to empower LLMs with causal-and-effect knowledge and address the causal reasoning challenges [17, 2, 16]. For instance, Jin et al. [17] introduce the CORR2CAUSE benchmark dataset on which they fine-tune their model. This is done to both asses and empower LLMs with causal reasoning ability. In fact, CORR2CAUSE dataset serves as a tool to evaluate the proficiency of LLMs in discerning causal relationships, particularly when the LLMs are fine-tuned to distinguish causation from correlational statements in the context of NLP.
  • Performance Evaluation: The second category focuses on using LLM for causal discovery and delves into emerging research that explores the causal analysis capabilities of Large Language Models. In contrast to causal discovery algorithms relying on statistical patterns in the data, this group utilizes LLMs to discover causal structures from variables. A majority of these methods solely utilize LLMs to predict pairwise causal relationships among a given set of variables [46, 24, 20, 41, 30, 6, 49].
  • Prior or Posterior Knowledge: In the third category, focused on employing LLMs, the objective is either to assign direction to undirected edges generated by causal discovery algorithms or to impose constraints on the edge orientation and functionality of these algorithms.
    [7, 6, 43].
Despite these efforts from conventional data-driven causal discovery algorithms to propose robust, precise, adaptable, efficient, and scalable causal discovery algorithms, encountered limitations and inefficiencies persist. These challenges are as follows. 1) Real-world data, often sparse and insufficient for accurately capture authentic probability distributions [6]. 2) Sole reliance on pre-collected static data introduces accuracy risks, particularly when models must adapt to dynamic real-world data and unforeseen factors. 3) Inferring complete edge orientations from observed data is hindered by the existence of equivalent Directed Acyclic Graphs (DAGs) [6, 35]. 4) Algorithm dependence on domain knowledge experts, who may be scarce, time/resource-intensive, or exhibit variable quality across domains [12]. 5) Traditional causal discovery algorithms fall short in answering user-submitted causal questions due to a lack of proficiency in language understanding and processing. These challenges collectively contribute to diminished accuracy, incompleteness, and unreliability in the estimated causal graph.
On the other hand, significant advances have been made in utilizing LLMs for causal tasks. However, their inherent limitations in precision and complexity handling remain evident. These challenges are highlighted as follows. 1) LLMs inherently lack the precision necessary for accurately responding to complex, user-generated causal queries [41]. 2) LLMs are limited in their ability to dissect and comprehend nuanced causal concepts without additional data-driven causal reasoning algorithms. 3) There is a challenge in constructing complete causal graphs and unraveling intricate causal relations due to the oversimplified understanding of LLMs. 4) LLMs struggle with handling extensive datasets, often failing to capture the depth and variability within them. These issues collectively hinder the effectiveness of LLMs in accurately and reliably determining causal relationships. Consequently, data-driven causal reasoning algorithms assume a critical role in mitigating the limitations of LLMs in causal tasks, offering nuanced comprehension of causal concepts, unraveling intricate causal relations, constructing complete causal graphs, and handling extensive datasets.
In light of these considerations, a unified, comprehensive causal framework that integrates LLMs with data-driven conventional causal discovery algorithms is required. To address this need, we propose the development of ALCM. ALCM aims to enhance the robustness and accuracy of causal discoveries by leveraging the conventional causal discovery algorithms and LLMs.
Table 1 indicates the capabilities of two distinct causal discovery methods—Conventional data-driven Causal Discovery (CCD), LLMs-based approaches, and ALCM framework—across essential functional attributes. Dynamic Data Adaptability[23, 5, 49] is the capability of a method to adjust to changing data, while Detection of Hidden Variables[23, 49] refers to identifying unobserved influencers within the dataset. Comprehensive Graph Model Representation [6] assesses the completeness of the depicted causal structure, and Predictive Accuracy[20, 39, 23, 43, 41, 30, 49] measures the success in forecasting the correct causal relations. CCD methods are limited by their reliance on pre-defined statistical models as well as domain knowledge expert validation, lacking adaptability to dynamic data, generalizability[20, 15] to unseen data, autonomy, and lack of accuracy. Similarly, while LLMs are adept at dynamicity of data, generalizability, and detecting hidden variables, they fall short in providing comprehensive graph model representations, interpretability, explainability, autonomy, and precision for causal discovery task. ALCM combining these strengths while enhancing user independence from expert validation [20] and interpretability[8] in causal discovery.
Table 1: Comparative Analysis of CCD, LLMs, and ALCM across Key Functional Attributes
Descriptive Attribute CCD 1 LLMs ALCM
Dynamic Data Adaptability $\times$ $\checkmark$ $\checkmark$
Detection of Hidden Variables $\times$ $\checkmark$ $\checkmark$
Comprehensive Graph Model Representation $\checkmark$ $\times$ $\checkmark$
Predictive Accuracy $\times$ $\times$ $\checkmark$
Autonomous Operation $\times$ $\times$ $\checkmark$
Generalizability to Unseen Data $\times$ $\checkmark$ $\checkmark$
Autonomous Expert Validation $\times$ $\checkmark$ $\checkmark$
Interpretability and Explainability $\checkmark$ $\times$ $\checkmark$
  • 1
    CCD methods often rely on pre-defined statistical models and assumptions about the data generation process.
  • 2
    LLMs-based methods may utilize vast amounts of data and natural language processing to infer causal relationships, potentially incorporating domain expertise.
  • 3
    ALCM synthesizes the strengths of both CCDs and LLMs to uncover causal connections.
3 Proposed Framework
In this section, we present ALCM, an advanced causal discovery framework aimed to leverage the combined strengths of traditional causal discovery algorithms and LLMs. ALCM provide an automated pipeline constructing a comprehensive causal graph, refining it, and incorporating previously overlooked insights to enrich the resulting causal model. This integration aims to utilize the precision of conventional causal discovery algorithms in identifying data relationships, while also enhancing and validating these findings with insights from LLMs. Fig. 1 indicates an overview of the ALCM framework. The algorithmic perspective of the ALCM framework is detailed in Algorithm 1. The ALCM framework includes three principal components: Causal Structure Learning, Causal Wrapper, and LLM-driven Refiner. To clarify the functionality and definitions of the framework, we present and exemplify these components in the following.
Figure 1: ALCM Architecture
Algorithm 1 ALCM
1:Observed dataset, $O$; Contextual Causal Information, $C$; Metadata, $M$
2:Causal DAG, $DAG$
3:Initialize and run the selected data-driven causal discovery algorithms CD, $G_{\text{i}}\leftarrow CD(O)$
4:Generate the causal prompt by injecting C and M
5:for each $z=(e_{i},e_{j})$ in $G_{\text{i}}$ do
6:     if $z$ is validated by LLM-Driven Refiner then
7:         $G_{\text{i}}\leftarrow G_{\text{i}}\cup\emptyset$
8:     end if
9:     if $z$ orientation is revised by LLM-Driven Refiner then
10:         $G_{\text{i}}\leftarrow z^{\prime}\cup G_{\text{i}}$
11:     end if
12:     if $z$ is removed by LLM-Driven Refiner then
13:         $G_{\text{i}}\leftarrow G_{\text{i}}-z^{\prime}$
14:     end if
15:     if a new $z^{\prime\prime}$ is added by LLM-Driven Refiner then
16:         $G_{\text{i}}\leftarrow z^{\prime\prime}\cup G_{\text{i}}$
17:     end if
18:end for
19:return $G_{\text{i}}$
3.1 Causal Structure Learning
The Causal Structure Learning component is our data-driven conventional causal discovery component. It receives a dataset as its input and generates the initial causal graphs from the dataset. This component uncovers causal insights by analyzing purely observational data, and it develops graphical structures that can be interpreted causally. This component directly influences the accuracy and reliability of both the final causal graph and future causal inferences drawn from the data.
Conventional causal discovery algorithms can be leveraged in this component. For instance, we can implement conventional causal discovery algorithms, including Peter-Clark (PC) [38] and Linear Non-Gaussian Acyclic Model (LiNGAM) [37] to discern the probabilistic dependencies and independencies among variables. These algorithms are selected based on their proven efficacy in uncovering intricate causal relationships within complex data.
For the implementation part, we implement PC conventional causal discovery algorithms; moreover, we propose a hybrid method combining the PC and LiNGAM algorithms to identify causal relationships effectively. Building on this foundation, the causal structure learning component crafts an initial causal graph, encapsulating the potential causal linkages derived from the datasets. Finally, this generated causal graph is then relayed to the Causal Wrapper component for further processing.
3.2 Causal Wrapper
The Causal Wrapper component serves as a critical intermediary or bridge between the causal structure learning and LLM-driven refiner components. This component encapsulates and translates the raw, initial causal graph into a series of contextual, causal-aware prompts (i.e., causal prompts). These prompts are fed to the LLM-driven refiner. The primary aim of these causal prompts is to act as guides for the LLM-driven refiner, aiding it in comprehending the initial causal graph. Furthermore, these causal prompts direct the LLM-driven refiner to identify and integrate the relevant and updated causal background knowledge to make the solution more suited to the specific causal discovery problem at hand. Given these reasons, this prompting strategy ensures that the final causal graph is not only precise but also robust and reflective of the underlying causal mechanisms within the dataset.
Equation 1 shows our causal-aware prompting strategy by infusing the context of problem and metadata information into the prompts. This prompting strategy was inspired from an effort by Kim et al. [21]. They demonstrated that contextual information is important in boosting the overall performance of LLMs’ responses.
[…]
This enhancement is accomplished by incorporating explicit elements into the prompt, with each edge being transformed into a causal prompt structured as follows:
Instructions: This section clarifies the role of LLMs, their objectives, and the expected behavior. Causal Context: It includes details about the selected causal discovery algorithm, such as its name and output. Metadata: This section outlines the dataset domain or variable names along with their descriptions. Question: It specifies the precise query, for example, whether A causes B. Output format: This delineates the desired format for the output.
Figure 2 illustrates an example of the causal wrapper’s functionality. Additionally, the output can incorporate supplementary reasoning and confidence levels for the generated response. For instance, a simple instruction can prompt the LLM-driven refiner to engage in step-by-step reasoning, employing a Chain-of-Thought (CoT) approach [45]. Moreover, it can request the LLM to indicate its level of confidence or likelihood regarding the generated output, using either a log-likelihood value or a confidence level. Once these causal prompts are generated, they are dispatched to the LLM-driven refiner component. This method ensures that the ALCM framework optimally utilizes LLMs for uncovering, refining, and validating causal relationships, thereby advancing the field of causal discovery with a high level of accuracy.
Figure 2: Causal Prompt Demonstration
3.3 LLM-driven Refiner
The LLM-driven Refiner leverages advanced language models in the refinement and evaluation of causal graphs. This component receives a series of intricately designed, contextual causal prompts from the causal wrapper component, which serve as a nuanced guide for its operations. Its core function is to assess, refine, and potentially augment the initial causal graph by evaluating the causal edges/nodes, and, where necessary, adding or removing nodes and edges to represent the underlying causal mechanisms better.
The significance of the LLM-driven Refiner lies in its capacity to address and alleviate inherent limitations present in both the causal discovery algorithms and the datasets themselves. This component plays a pivotal role in uncovering and assimilating previously overlooked or concealed causal information, thereby elevating the accuracy and comprehensiveness of the causal graph. The identification and integration of hidden causal relationships into the graph are essential, as they can reveal causal connections or nodes that traditional causal discovery methods might miss or that dataset constraints could obscure. Upon completion of the refinement process, the results are saved, and various post-processing techniques are applied to generate the final graph. These techniques involve leveraging natural language processing (NLP) to parse and extract causal relationships from textual responses provided by LLMs. Subsequently, these extracted relationships undergo validation and structuring to form a coherent causal graph.
4 Implementation
We elucidate the technical underpinnings and strategic choices behind the deployment of the ALCM framework. We provide two demonstrations of implementation of our framework to show that our framework can enhance the accuracy and generalizability.
4.1 Implementation 1 (ALCM-PC)
For the first implementation, we select PC causal discovery algorithm. The PC algorithm is renowned for its robustness in dealing with large datasets and its ability to infer causal structures through conditional independence tests, making it highly efficient in uncovering complex causal networks. For the causal wrapper component, we utilize causal prompt. We illustrate one example of our prompt in Figure 3. For LLM-driven refiner, we exploit OpenAI GPT-4 [29, 3] in our pipeline.
Figure 3: Prompt Template
4.2 Implementation 2 (ALCM-Hybrid)
For the second implementation, we leverage a hybrid approach (which outlined in Section 3.1) including PC [38] and LiNGAM [37] algorithms due to their complementary strengths and proven effectiveness in identifying causal relationships. This hybrid method utilizes a majority vote mechanism for identifying common causal edges recognized by both algorithms. For edges that are uniquely identified by only one algorithm and not common to both, we introduce an extra step by employing LLMs as a decisive judge. This entails presenting these edges to LLMs to ascertain their potential as causal links based on contextual understanding and causal reasoning capabilities. If affirmed, these edges are added to an augmented set of causal connections. The causal wrapper component applies the causal prompt template and the result is sent to the LLM-driven refiner. The LLM-driven refiner (OpenAI model) Component evaluates, refines, and enhances the causal graph to produce a final, enhanced causal structure.
5 Experiments
In this section, we first present benchmark datasets used in our expermients. Next, we outline the evaluation metrics selected to measure the framework’s performance in terms of accuracy, robustness, and reliability. Finally, we summarize the experimental results, demonstrating the effectiveness of the ALCM framework in generating and refining causal graphs, and its ability to reveal latent causal relationships, showcasing its advancement over existing methods.
5.1 Benchmark Datasets
We utilize six benchmark datasets and their ground truth causal graphs from the BN repository: Asia, Cancer, Child, Insurance, Sachs, Sangiovese [36, 25], and also the well-known Neuropathetic dataset [42] to evaluate the efficacy of the ALCM framework. These datasets are chosen for their diverse origins and complexities, covering a range of scenarios from medical studies to insurance modeling and genetic pathways. The importance of utilizing these benchmark datasets lies in their ability to provide a standardized basis for comparison, enabling the assessment of the ALCM framework’s performance across varied domains and conditions. Table 2 indicates a summary of these datasets.
Table 2: Summary of Datasets
Dataset Domain #Nodes #Edges
Asia Social Science 8 8
Cancer Medical 11 18
Child Social Science 20 31
Insurance Finance 27 43
Sachs Biological 11 18
Neuropathic Medical 221 475
Sangiovese Social Science 36 47
To ensure these datasets are compatible with the input requirements of causal discovery algorithms within the ALCM framework, we implement a series of preprocessing techniques as part of the causal structure learning component. This preprocessing involves cleaning the data, handling missing values, and normalizing data formats, among other adjustments, to tailor the datasets for optimal processing. By meticulously preparing these datasets, we facilitate their effective use as inputs for the causal discovery algorithms, ensuring that the initial causal graphs generated are as accurate and informative as possible.
5.2 Evaluation Metrics
We select five metrics to assess the effectiveness and precision of the ALCM framework’s causal discovery capabilities. The evaluation of the predicted causal graphs against the ground truth is paramount to validate the accuracy and reliability of our methodology. To this end, we employ five key metrics: precision, recall, F1-score, accuracy, and Normalized Hamming Distance (NHD), each selected for its ability to provide a comprehensive understanding of the framework’s performance from different perspectives [48].
  • Precision: measures the proportion of correctly identified causal relationships out of all relationships identified by the algorithm. This metric is crucial for ensuring that the causal links proposed by our framework are indeed valid, minimizing false positives.
  • Recall: assesses the fraction of true causal relationships that have been correctly identified by the algorithm, highlighting the framework’s ability to uncover the full extent of causal connections present within the data.
  • F1-score: serves as a harmonic mean of precision and recall, offering a single metric that balances both the accuracy and completeness of the identified causal relationships. This is particularly useful for comparing the overall performance of different causal discovery approaches.
  • Accuracy: evaluates the overall correctness of the causal graph, including both the presence of true causal connections and the absence of false ones. This metric provides a straightforward assessment of the model’s overall predictive performance.
  • Normalized Hamming Distance (NHD): quantifies the difference between the predicted causal graph and the ground truth by measuring the proportion of mismatched edges, adjusted for the size of the graph. NHD is instrumental in assessing the structural similarity of the causal graphs, offering insights into the nuanced differences that may not be captured by other metrics. In the context of a graph with m nodes, the NHD between the predicted graph G p and the ground-truth graph G is determined by calculating the number of edges that exist in one graph but not the other. This count is then divided by the total number of all possible edges–this formula is defined in Equation 2. In essence, the NHD provides a normalized measure of dissimilarity, offering insights into the accuracy of the predicted graph compared to the ground-truth graph, accounting for the total potential edges in the graph with m nodes.
$NHD=\sum_{i=1}^{m}\sum_{j=1}^{m}\frac{1}{m^{2}}\cdot 1,\quad\text{where }G_{ij}\neq G_{p_{ij}}.$ (2)
5.3 Experimental Results
In this section, we present our experimental results and a comprehensive analysis of the performance of the ALCM framework, utilizing the seven aforementioned datasets and the five evaluation metrics to benchmark against existing methodologies in causal discovery, including conventional algorithms and approaches leveraging LLMs. Specifically, we implement the PC algorithm as a representative of conventional causal discovery methods and utilize LLMs powered by OpenAI’s technology as a cutting-edge counterpart. LLM-based approaches generates the pairwise sets of nodes and analyze them. The outcomes of these evaluations are indicated in Table 3, which underscores the substantial enhancements achieved by our ALCM framework across various metrics. We also include the implementation of PC as the backbone of causal discovery algorithm in ALCM-PC and for ALCM-Hybrid, we employ the hybrid algorithm.
Table 3: Evaluation Results for Various Causal Discovery Methods
Dataset Method Metrics
Precision Recall F1-Score Accuracy NHD
Asia PC 0.75 0.375 0.5 33.33 0.1429
LLMs 0.1428 0.2174 0.1742 16.00 0.75
ALCM-PC 1.0 0.5945 0.746 87 0.0893
ALCM-Hybrid 0.8889 1.0 0.9412 96.55 0.0179
Cancer PC 0.5 0.5 0.5 33.33 0.2
LLMs 0.158 0.75 0.261 21.4 0.85
ALCM-PC 0.667 1.0 0.800 85.71 0.1
ALCM-Hybrid 0.9333 0.9655 0.9492 90.32 0.0333
Child PC 0.20 0.28 0.233 27.00 0.121
LLMs 0.0657 0.48 0.1156 29.21 0.8765
ALCM-PC 1.0 0.6185 0.764 78.89 0.047
ALCM-Hybrid 0.95 0.72 0.819 98.00 0.018
Insurance PC 0.2153 0.2692 0.2393 13.59 0.8640
LLMs 0.069 0.5833 0.1234 22.90 0.8620
ALCM-PC 1.0 0.857 0.923 94.8 0.054
ALCM-Hybrid 1.0 0.9 0.947 96.4 0.037
Neuropathetic PC 0.45 0.551 0.4954 51.7 0.1364
LLMs 0.105 0.2831 0.202 10.2 0.4537
ALCM-PC 0.8846 0.6201 0.7291 89.26 0.0575
ALCM-Hybrid 0.8667 0.9692 0.9151 98.00 0.0165
Sachs PC 0.4167 0.5882 0.4878 80.91 0.209
LLMs 0.2081 0.6471 0.3149 63.24 0.9051
ALCM-PC 0.6117 0.7059 0.6554 87.5 0.1881
ALCM-Hybrid 0.7294 1.0 0.8435 90.44 0.1727
Sangiovese PC 0.4348 0.1818 0.2564 14.71 0.2761
LLMs 0.2880 0.6545 0.400 25.0 0.5143
ALCM-PC 0.6548 1.0 0.7914 65.48 0.1381
ALCM-Hybrid 0.8209 1.0 0.9016 93.41 0.0659
Our analysis shows that, compared to the baseline PC algorithm and LLM-based approaches, both ALCM-PC and ALCM-Hybrid demonstrate a notable increase in precision, recall, F1-score, and accuracy, indicating a significant improvement in both the reliability and completeness of the identified causal relationships. Conversely, the NHD exhibits a marked decrease for both ALCM-PC and ALCM-Hybrid suggesting a closer structural alignment with the ground truth causal graph and, therefore, a more accurate representation of the causal dynamics within the datasets. The superior performance of ALCM-Hybrid over ALCM-PC can primarily be attributed to its use of a dual strategy that combines conventional causal discovery algorithms and employs a majority voting mechanism, alongside leveraging LLMs to incorporate the latest information from the Internet. ALCM-PC and ALCM-Hybrid’s superior performance stem from blending conventional causal discovery techniques with LLMs and an automated refinement pipeline. This innovative mix not only utilizes conventional methods for identifying causal links but also benefits from LLMs’ ability to process contextual information and updates. This synergy significantly enhances causal relationship accuracy and graph reliability. The result is a robust causal discovery tool that demonstrates marked improvements in key metrics and a closer alignment with the ground truth causal dynamics. We can also observe a low precision and high NHD values for some results of LLMs-based approaches, suggesting that a significant portion of the relationships identified by the LLMs were not actually present in the ground truth, highlighting a potential issue with the algorithm’s specificity or its tendency to overgeneralize from the input data.
We depict the causal graphs obtained by a couple of causal discovery methods on Sachs dataset in Figure 4. The Sachs dataset [36] includes data on 11 phosphorylated proteins and phospholipids from human immune cells, providing a basis for analyzing protein signaling pathways and constructing causal networks. It is especially valuable for causal discovery research, with data collected from cells under different experimental conditions, making it an excellent benchmark for testing causal discovery algorithms. Graph of ground truth, LLMs-based approach, PC, ALCM, ALCM-Hybrid are shown in Figures 4(a), 4(b), 4(c), 4(d), 4(e), respectively.
(a) Causal graph for Ground Truth
(b) Causal graph for LLMs-based
(c) Causal graph for PC
(d) Causal graph for ALCM
(e) Causal graph for ALCM-Hybrid
Figure 4: Causal graphs Demonstrations
The enhanced performance across all metrics for both ALCM-PC and ALCM-Hybrid variants can be directly linked to their innovative methodologies. ALCM’s use of LLMs introduces a layer of causal reasoning and validation that is absent in traditional approaches, while the hybrid model further capitalizes on this by combining algorithmic precision with AI’s contextual insights. This strategic amalgamation ensures that our framework is at the forefront of causal discovery, setting a new benchmark for accuracy, comprehensiveness, and applicability in the field. We also visualize the additive contributions of each causal discovery framework in Figures 5 and 6 on two benchmarks–neuropathetic and sachs.
Figure 5: Additive Contribution on Causal Discovery Accuracy on Neuropathetic Pain
Figure 6: Additive Contribution on Causal Discovery Accuracy on Sachs
5.4 Results for Adding New Nodes or Edges
The extensive updated knowledge and expert supervision provided by LLMs significantly facilitate the identification of elusive variables (Markov blanket) and causal connections. These might remain undetected or in the dataset or overlooked by causal discovery algorithms. Figure7(a) and 7(b) show these capability of unmasking these hidden aspects. As Figure 7(a) demonstrates, the causal discovery algorithm (PC) fails to detect all of the true nodes and edges, but ALCM can provide us with new nodes or edges that not present in the output set of causal discovery algorithm as illustrated in 7(b). We prompted LLMs to provide us the confidence level for its responses as well. The validity of ALCM answer is also confirmed by the up-to-the-date medical articles, including [28].
(a) Causal graph for PC
(b) Causal graph for ALCM
Figure 7: Causal graphs for demonstrating new nodes or edges
The traditional causal discovery depends on the structured dataset and their quality which are curated and annotated by human experts. However, these dataset are neither available in a wide range of domains or can be generalize to the new tasks. Hence, we empower ALCM by viture of LLMs component with this capability to uncover hidden variables and causal connections. Figure 8 indicates the ALCM capability to entangle the hidden variables and causal relations which are not present in the dataset.
Figure 8: Results for Uncover, Hidden, or Ignored Nodes and Edges
6 Future Work
In the subsequent phases of our research, we aim to develop a more sophisticated causal-aware framework. This framework will leverage the power of knowledge graphs, which are instrumental in augmenting the accuracy of our models. Furthermore, we plan to explore the integration of our framework with Monte Carlo Tree Search (MCTS). This integration is envisioned to evolve our system into a more dynamic and adaptive problem-solving agent.
Additionally, to advance the ALCM framework’s capabilities and address the issue of LLM hallucination, we propose integrating ALCM with the Retrieval-Augmented Generation (RAG) system and openCHA [1]. This integration aims to harness RAG’s ability to augment LLMs’ generative processes with data retrieval, ensuring that causal discovery are grounded in relevant and factual information. openCHA sophisticated dialogue capabilities will further enhance ALCM by enabling dynamic, interactive validation of causal hypotheses.
7 Conclusion
This study underscored the transformative potential of combining LLMs with data-driven causal discovery algorithms through the introduction of the Autonomous LLM-Augmented Causal Discovery Framework (ALCM). The ALCM emerges as a groundbreaking solution, aiming to enhance the generation of causal graphs by leveraging the sophisticated capabilities of LLMs alongside conventional causal discovery techniques. By integrating causal structure learning, a causal wrapper, and an LLM-driven causal refiner, ALCM facilitated an autonomous approach to causal discovery, significantly outperforming existing methodologies in both accuracy and interpretability. The empirical validation of ALCM not only attests to its superior efficacy over prevailing LLM methods and conventional causal reasoning mechanisms but also illuminates new pathways for leveraging LLMs in uncovering intricate causal relationships across a myriad of domains.